the no. of moles of hydroxide ions in 0.3litre of 0.005M solution of Ba(OH)2  is

Grade:12th pass

3 Answers

Ashish Kumar
7 years ago
By the formula- Molarity=Moles of solute/Volume of solution we get the moles of solute is 0.0013 moles. Now the volume of water is 0.3L. The volume of 0.0013 moles of Ba(OH)2 equals to 0.0013 x 22.4L= 0.029L. Therefore the volume os solvent i.e. water = 0.3L- 0.029L=0.271L. Now 22.4L is the volume of 1 mole of water so by unitary method we get the volume of 0.271L of water= 0.271/22.4=0.0121 moles.
Now moles of Ba(OH)2 =0.0013. In 1 molecule of the compound there are 2 molecules of hrdroxide ions so the number of moles of hydroxide ions= 0.0013 x 2 =0.0026.
aMoles of water =0.0121 moles. 1 molecule of water has 1 H+ ion and 1 OH- ion so moles of OHions gained from water = 0.0121 x 1 =0.0121 moles.
Therefore total moles of hydroxide ions = 0.0026+0.0121=0.0147 Moles
Hithesh seedarla
7 years ago
Ba(OH)2=0.005M, in 0.03litre no.of moles of (OH-)=?       
sol: M=n/v  in litres 
0.05 =n/o.3
0.05×0.3=n
n(OH-)=0.0015
Mohammed Mushthaq
6 years ago
1M of a solution means , 1L of the solution contain 1 mole of solute.Hence the given solution contains 0.005 moles of Ba (OH)2 in 1L. So,there will be 0.0015 (0.005×0.3) moles of Ba (OH)2 in 0.3 L of the solution.1 mol Ba (OH)2 contains 2 mol of OH-.Hence,0.0015 mole contains 0.003 moles of OH- ion. Hope this helps.